### v4

# Inverse scattering for $N$-body time-decaying harmonic oscillators

Abstract
In the previous study [26], the author proved the uniqueness of short-range potential functions using the Enss-Weder time-dependent method [16] for a two-body quantum system described by time-decaying harmonic oscillators. In this study, we extend the result of [26] to the $N$-body case. We use the approaches developed in [16, 49, 53] to prove that the high-velocity limit of the scattering operator uniquely determines all the pairwise interaction potentials among the NN particles, focusing respectively on each fixed pair of particles.
Keywords: NN-body scattering theory, wave operator, scattering operator
MSC2020: 35R30, 81U10, 81U40
1 Introduction
We consider a quantum system governed by the $N$-particle harmonic oscillators and interactional potential functions described by the Hamiltonian
$\tilde{H}(t)=-\sum_{j=1}^{N}\Delta_{r_{j}}/(2m_{j})+k(t)\sum_{j=1}^{N}m_{j}r_{j}^{2}/2+\sum_{1\leqslant j<k\leqslant N}V_{jk}(r_{k}-r_{j})$ (1.1)
acting on $L^{2}({\mathbb{R}^{dN}})$ with $N\geqslant 2$ and $d\geqslant 2$, where $r_{j}\in\mathbb{R}^{d}$ is the position of the $j$th particle, $m_{j}>0$ is the mass, $\Delta_{r_{j}}$ is the Laplacian for $r_{j}$, and $V_{jk}$ is the interaction between $j$th and $k$th particles. The time-dependent coefficient $k(t)$ is defined as
$k(t)=\begin{cases}\ \omega^{2}&\quad\mbox{if}\quad|t|<T,\\ \ \sigma/t^{2}&\quad\mbox{if}\quad|t|\geqslant T\end{cases}$ (1.2)
for $0<\sigma<1/4$, $\omega>0$ and $T>0$.
We now introduce the Jacobi coordinates in which the center of mass of all $N$ particles is located at the origin (see [31, Chapter 3] or [45, XI.5]). Let $\mathscr{X}$ be the subspace of $\mathbb{R}^{dN}$ described by
$\mathscr{X}=\left\{(r_{1},\ldots,r_{N})\in\mathbb{R}^{dN}\Biggm|\sum_{j=1}^{N}m_{j}r_{j}=0\right\}\simeq\mathbb{R}^{d(N-1)}$ (1.3)
and $\mathscr{X}_{\rm cm}$ be the orthogonal space for $\mathscr{X}$ associated with the scalar product
$(r,\tilde{r})_{\mathbb{R}^{dN}}=\sum_{j=1}^{N}m_{j}r_{j}\cdot\tilde{r}_{j}$ (1.4)
on $\mathbb{R}^{dN}$ for $r=(r_{1},\ldots,r_{N})$ and $\tilde{r}=(\tilde{r}_{1},\ldots,\tilde{r}_{N})\in\mathbb{R}^{dN}$. When the position of the center of mass is removed, the unperturbed Hamiltonian of (1.1) is given by
$H_{0}(t)=-\sum_{j=1}^{N}\Delta_{r_{j}}/(2m_{j})+\Delta_{\rm cm}/(2M)+k(t)\sum_{j=1}^{N}m_{j}r_{j}^{2}/2-k(t)Mr_{\rm cm}^{2}/2$ (1.5)
acting on $L^{2}(\mathscr{X})$, where $M=\sum_{j=1}^{n}m_{j}$ is the total mass, $r_{\rm cm}=\sum_{j=1}^{N}m_{j}r_{j}/M\in\mathscr{X}_{\rm cm}$ is the position of the center of mass, and $\Delta_{\rm cm}$ is the Laplace-Beltrami operator on $\mathscr{X}_{\rm cm}$. This expression (1.5) can be simplified by rewriting it in Jacobi coordinates as
$y_{j}=r_{j+1}-\sum_{k=1}^{j}m_{k}r_{k}/\sum_{k=1}^{j}m_{k}\in\mathscr{X}$ (1.6)
for $1\leqslant j\leqslant N-1$. The reduced mass $\mu_{j}$ is defined as
$1/\mu_{j}=1/\sum_{k=1}^{j}m_{k}+1/m_{j+1}.$ (1.7)
Under these notations, we have
$\sum_{j=1}^{N}m_{j}r_{j}^{2}=Mr_{\rm cm}^{2}+\sum_{j=1}^{N-1}\mu_{j}y_{j}^{2},$ (1.8)
$\sum_{j=1}^{N}\Delta_{r_{j}}/m_{j}=\Delta_{\rm cm}/M+\sum_{j=1}^{N-1}\Delta_{y_{j}}/\mu_{j}$ (1.9)
where $\Delta_{y_{j}}$ is the Laplace-Beltrami operator on $\mathscr{X}$. Therefore, in these coordinates, (1.5) can be written as the unperturbed Hamiltonian
$H_{0}(t)=-\sum_{j=1}^{N-1}\Delta_{x_{j}}/2+k(t)\sum_{j=1}^{N-1}x_{j}^{2}/2$ (1.10)
acting on $L^{2}(\mathscr{X})$ with $x_{j}=\sqrt{\mu_{j}}y_{j}$ for simplicity.
We now give the assumptions for each potential function $V_{jk}$ in (1.1) as a perturbation of $H_{0}(t)$. Let $\lambda$ satisfy
$0<\lambda=(1-\sqrt{1-4\sigma})/2<1/2.$ (1.11)
This $\lambda$ describes the effective scaling rate induced by the time-decaying harmonic potential and is directly related to the classical trajectories $x(t)=O(|t|^{1-\lambda})$ as $|t|\rightarrow\infty$. The condition $0<\lambda<1/2$ ensures the short-range character of the interaction under the scaling associated with the propagator. The bracket $\langle\cdot\rangle$ is defined by $\sqrt{1+|\cdot|}$ and $A\lesssim B$ means that there exists a constant $C>0$ such that $A\leqslant CB$.
Assumption 1.1.
Each function $V_{jk}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ for $1\leqslant j<k\leqslant N$ is a multiplication operator and decomposed into two parts:
$V_{jk}=V_{jk}^{\rm bdd}+V_{jk}^{\rm sing}.$ (1.12)
The bounded part $V_{jk}^{\rm bdd}\in L^{\infty}(\mathbb{R}^{d})$ satisfies
$|V_{jk}^{\rm bdd}(x)|\lesssim\langle x\rangle^{-\rho}$ (1.13)
with $\rho>1/(1-\lambda)$. The singular part $V_{jk}^{\rm sing}\in L^{q}(\mathbb{R}^{d})$ is compactly supported on $\mathscr{X}$, where the Lebesgue exponent $q$ satisfies
$\infty>q\begin{cases}\ =2\quad&\mbox{\rm if}\quad n\leqslant 3,\\ \ >n/2\quad&\mbox{\rm if}\quad n\geqslant 4.\end{cases}$ (1.14)
It is well-known that each $V_{jk}^{\rm sing}$ is $\Delta-$bounded infinitesimally. The interacting Hamiltonian is
$H(t)=H_{0}(t)+\sum_{1\leqslant j<k\leqslant N}V_{jk}(r_{k}-r_{j})$ (1.15)
acting on $L^{2}(\mathscr{X})$. Here we note $r_{k}-r_{j}\in\mathscr{X}$ for any $1\leqslant j<k\leqslant N$. Under these definitions, the perturbed full Hamiltonian (1.1) is decomposed into
$\tilde{H}(t)=H(t)\otimes 1+1\otimes(-\Delta_{\rm cm}/(2M)-k(t)Mx_{\rm cm}^{2}/2)$ (1.16)
on $L^{2}(\mathscr{X})\otimes L^{2}({\mathscr{X}_{\rm cm}})$, where $1$ denotes the identities on $L^{2}(\mathscr{X})$ and $L^{2}(\mathscr{X}_{\rm cm})$. Therefore, we will concentrate on studying the quantum system governed by $H(t)$. Because two parameter unitary propagators generated by $H_{0}(t)$ and $H(t)$ exist uniquely by the result of [54, Theorem 6 and Remark (a)], we denote these propagators as $U_{0}(t,s)$ and $U(t,s)$. In $N$-body quantum scattering theory, we have to consider the cluster decomposition $a=\{C_{1},\ldots,C_{m}\}$ with $m\geqslant 2$, where $C_{j}\subset\{1,\ldots,N\}$ for $1\leqslant j\leqslant m$, which satisfies $\cup_{j=1}^{m}C_{j}=\{1,\ldots,N\}$ and $C_{j}\cap C_{k}=\emptyset$ if $j\not=k$. Let $\mathscr{X}^{a}$ be the subspace of $\mathscr{X}$ given by
$\mathscr{X}^{a}=\left\{(r_{1},\ldots,r_{N})\in\mathscr{X}\Biggm|\sum_{j\in C}m_{j}r_{j}=0,\ C\in a\right\}$ (1.17)
and $\mathscr{X}_{a}$ be its orthogonal space. The cluster Hamiltonian is defined as
$H_{a}(t)=-\sum_{j=1}^{N-1}\Delta_{x_{j}}/2+k(t)\sum_{j=1}^{N-1}x_{j}^{2}/2+\sum_{\{j,k\}\subset a}V_{jk}(r_{k}-r_{j})$ (1.18)
acting on $L^{2}(\mathscr{X})$. In the subspace $\mathscr{X}_{a}$, $V_{jk}$ does not vanish for $\{j,k\}\subset a$. Therefore, it is natural that the cluster interaction
$\sum_{\{j,k\}\not\subset a}V_{jk}(r_{k}-r_{j})$ (1.19)
be regarded as the perturbation and that the cluster Hamiltonian $H_{a}(t)$ of (1.18) be considered the unperturbed system for any cluster decomposition $a$. However, to discuss inverse scattering, it suffices to consider the $N$-cluster case where $a=\{\{1\},\ldots,\{N\}\}$ and the wave operators defined as
$W^{\pm}=\operatorname*{s-lim}_{t\rightarrow\infty}U(t,0)^{*}U_{0}(t,0).$ (1.20)
We give the proof of existence of (1.20) in Section 2. Then the scattering operator is defined as
$S(V)=(W^{+})^{*}W^{-}$ (1.21)
for $V=\sum_{1\leqslant j<k\leqslant N}V_{jk}(r_{k}-r_{j})$.
Theorem 1.2.
Let $V_{1}=\sum_{1\leqslant j<k\leqslant N}V_{jk,1}$ and $V_{2}=\sum_{1\leqslant j<k\leqslant N}V_{jk,2}$ satisfy Assumption 1.1. If $S(V_{1})=S(V_{2})$, then $V_{jk,1}=V_{jk,2}$ for $1\leqslant j<k\leqslant N$.
Many researchers have studied scattering theory under the $N$-body quantum systems, in particular, the the existence and asymptotic completeness of the wave operators, has been studied. Regarding the standard $N$-body Schrödinger operators, see [5, 13, 14, 18, 46, 47, 48]; the detailed history and proofs are also provided in [15, 31]. For $N$-body Schrödinger operators in the external electric fields, see [1, 3, 8, 9, 20]. For the external magnetic fields, see [2, 17]. However, there are no results regarding scattering theory for $N$-body time-decaying harmonic oscillators.
The Enss-Weder time-dependent method was invented in [16] for the standard Schödinger operators. Since then, many researchers have applied this method to other quantum models. The Stark effect and time-dependent electric fields were studied by [4, 6, 7, 10, 11, 22, 25, 42, 44, 49]. The repulsive Hamiltonians were studied by [21, 24, 44]. The fractional Laplacian and Dirac equations were studied by [23, 32]. The non-linear Schrödinger equations were studied by [50]. The Hartree-Fock equations were studied by [51, 52]. Recently, the author investigated the two-body time-decaying harmonic oscillator and repulsive Hamiltonian in [26, 27]. $N$-body inverse scattering was investigated by [16] for the standard Schrödinger case, and by [53, 49] for the Stark effect. Our proof also follows the strategies of these previous studies. That is, by an appropriate choice of Jacobi coordinates, the problem is reduced to the two-body case [26]. It should be emphasized that the pioneering works of [16, 53] already treated both the two-body and $N$-body case in a unified manner. The arguments of [16, 53, 49] show that the reduction to two-body channels is a fundamental principle.
Regarding scattering theory for the two-body time-decaying harmonic oscillator, [28] proved the existence of the wave operators under the assumption 1.1 only for the bounded part $V^{\rm bdd}$ and non-existence for $\rho\leqslant 1/(1-\lambda)$, and clarified that the condition $\rho>1/(1-\lambda)$ is short-range. We can intuitively understand the threshold $1/(1-\lambda)$ from the classical motion of the particle $x(t)=c_{1}t^{1-\lambda}+c_{2}t^{\lambda}$ for $t\geqslant T$, which satisfies the Newton equation $({\rm d}^{2}/{\rm d}t^{2})x(t)=-k(t)x(t)$ because $\lambda$ is one of the roots of the quadratic equation $\lambda^{2}-\lambda+\sigma=0$. The critical case $\sigma=1/4$ was studied by [29]. For the inverse square potential, [30] proved the asymptotic completeness of the wave operators. Moreover, [30] proved the Strichartz estimates and applied them to the initial-value problems for the non-linear Schrödinger equations. The non-linear Schrödinger equations and Strichartz estimates with time-decaying harmonic oscillators have also been studied by [33, 34, 35, 36, 37, 38, 39, 40].
In contrast to the time-independent harmonic case or standard free Schrödinger dynamics, the time-decaying quadratic potential induces a nontrivial scaling behavior in the classical trajectories and the quantum propagator. One of the key technical ingredients of this paper is a propagation estimate for the free dynamics associated with the time-dependent harmonic oscillator in the $N$-body framework (Proposition 2.2). Although the estimate itself can be reduced to the two-body case by an appropriate choice of Jacobi coordinates, its formulation in a form suitable for the $N$-body setting is essential for proving the existence of the $N$-cluster wave operators. Once the existence of the wave operators is established, the inverse scattering argument follows the Enss-Weder time-dependent method and reduces the problem to suitable two-body channels. While the inverse problem itself is treated by adapting known high-velocity arguments, the derivation of the propagation estimates (Lemmas 3.2 and 3.3) in the presence of a time-decaying harmonic potential and their extension from the two-body case studied in [26] to the $N$-body setting constitute the main technical novelty of the present work.
Throughout this paper, we identify $L^{2}(\mathscr{X})$ with $L^{2}(\mathbb{R}^{d(N-1)})$ and use the following notations. $\|\cdot\|$ denotes the $L^{2}$-norm or operator norm on $L^{2}(\mathbb{R}^{d(N-1)})$, and $(\cdot,\cdot)$ denotes the scalar product of $L^{2}(\mathbb{R}^{d(N-1)})$. $F(\cdots)$ is the characteristic function of the set $\{\cdots\}$.
2 Existence of Wave Operators
In this section, we prove the existence of the $N$-cluster wave operators (1.20). To do that, we reduce the problem to a much simpler form of strong limits. We define $D=(D_{1},\ldots,D_{N-1})$ for $D_{j}=-{\rm i}\nabla_{x_{j}}$ and $x=(x_{1},\ldots,x_{N-1})\in\mathbb{R}^{d(N-1)}$. Then the free Hamiltonian is written as
$H_{0}(t)=D^{2}/2+k(t)x^{2}/2=\sum_{j=1}^{N-1}D_{j}^{2}/2+k(t)\sum_{j=1}^{N-1}x_{j}^{2}/2.$ (2.1)
The propagators $U_{0}(t,s)$ and $U(t,s)$ have unitary factorizations for $t,s\geqslant T$ or $t,s\leqslant-T$ that were proved by [28, Proposition 1] in the two-body case. We define
$U_{0,\lambda}(t)=e^{{\rm i}\lambda x^{2}/(2t)}e^{-{\rm i}\lambda\log tA}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}$ (2.2)
if $t\geqslant T$ and
$U_{0,\lambda}(t)=e^{{\rm i}\lambda x^{2}/(2t)}e^{-{\rm i}\lambda\log(-t)A}e^{{\rm i}(-t)^{1-2\lambda}D^{2}/(2(1-2\lambda))}$ (2.3)
if $t\leqslant-T$, where $A=(D\cdot x+x\cdot D)/2$. Then the factorizations
$U_{0}(t,s)=U_{0,\lambda}(t)U_{0,\lambda}(s)^{*}$ (2.4)
and
$U(t,s)=e^{{\rm i}\lambda x^{2}/(2t)}e^{-{\rm i}\lambda\log|t|A}U_{\lambda}(t,s)e^{{\rm i}\lambda\log|s|A}e^{-{\rm i}\lambda x^{2}/(2s)}$ (2.5)
hold for $t,s\geqslant T$ or $t,s\leqslant-T$, where $U_{\lambda}(t,s)$ is the propagator generated by
$D^{2}/(2|t|^{2\lambda})+\sum_{1\leqslant j<k\leqslant N}V_{jk}(|t|^{\lambda}(r_{k}-r_{j})).$ (2.6)
Regarding the proofs of (2.4) and (2.5), it suffices to prove the following Proposition 2.1 because $e^{{\rm i}\lambda\log|t|A}D_{j}e^{-{\rm i}\lambda\log|t|A}=D_{j}/|t|^{\lambda}$ for $1\leqslant j\leqslant N$ clearly holds, and the rest of the proofs are demonstrated in [28, Proposition 1].
Proposition 2.1.
$e^{{\rm i}\lambda\log|t|A}(r_{k}-r_{j})e^{-{\rm i}\lambda\log|t|A}=|t|^{\lambda}(r_{k}-r_{j})$ (2.7)
holds for any $1\leqslant j<k\leqslant N$.
Proof of Proposition 2.1.
For $1\leqslant j<k\leqslant N$ fixed, we construct the Jacobi coordinates from $y_{1}=r_{k}-r_{j}$ and define $A_{1}=(x_{1}\cdot D_{1}+D_{1}\cdot x_{1})/2$ for $x_{1}=\sqrt{\mu_{1}}y_{1}$. We have
$e^{{\rm i}\lambda\log|t|A}y_{1}e^{-{\rm i}\lambda\log|t|A}=e^{{\rm i}\lambda\log|t|A_{1}}x_{1}e^{-{\rm i}\lambda\log|t|A_{1}}/\sqrt{\mu_{1}}.$ (2.8)
As in the proof of [28, Proposition 1], we have (2.7). ∎
To prove the existence of the wave operators (1.20), it suffices to prove the existence of the strong limits
$\operatorname*{s-lim}_{t\rightarrow\pm\infty}U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}$ (2.9)
by virtue of the factorizations (2.4) and (2.5), and the chain rules of the propagators. The following propagation estimate and the existence of the wave operators for the two-body time-decaying harmonic oscillator with only the bounded part $V^{\rm bdd}$ only were proved in [28, Proposition 1 and Theorem 1] by establishing the correspondence with (2.9) for the two-body case. Our proof is one of the extensions to the $N$-body case with the singular part $V_{jk}^{\rm sing}$.
Here we introduce a conical region in which all relative momenta are uniformly bounded away from zero. We denote a d(N1)d(N-1)-dimensional open ball with radius $r>0$ centered at $\xi\in\mathbb{R}^{d(N-1)}$ as $B_{d(N-1)}(\xi,r)$ and its spherical surface as $\mathbb{S}^{d(N-1)-1}$. For any $1\leqslant j<k\leqslant N$, there exists a full-rank matrix $L_{jk}:\mathbb{R}^{d(N-1)}\rightarrow\mathbb{R}^{d}$ such that any conjugate momentum $\xi_{jk}\in\mathbb{R}^{d}$ associated with the relative coordinate $r_{k}-r_{j}$ is given by $\xi_{jk}=L_{jk}\xi$ for some $\xi\in\mathbb{R}^{d(N-1)}$. We note that $L_{jk}$ is surjective and that $\dim\ker L_{jk}=d(N-2)$. Let $\omega_{0}\in\mathbb{S}^{d(N-1)-1}\setminus(\cup_{\leqslant j<k\leqslant N}\ker L_{jk})$ be fixed. Then $|L_{jk}\omega_{0}|>0$ holds. In particular, there exist $\epsilon_{0}>0$ and $\delta>0$ such that $|L_{jk}\omega|\geqslant\epsilon_{0}$ for any $\omega\in\mathbb{S}^{d(N-1)-1}\cap B_{d(N-1)}(\omega_{0},\delta)$. We then define a conical region
$\mathscr{C}_{\omega_{0},\delta}=\{\xi\in\mathbb{R}^{d(N-1)}\setminus\{0\}\bigm|\xi/|\xi|\in\mathbb{S}^{d(N-1)-1}\cap B_{d(N-1)}(\omega_{0},\delta)\}.$ (2.10)
By this definition, if $\xi\in\mathscr{C}_{\omega_{0},\delta}$, then $|\xi_{jk}|=|L_{jk}\xi|\geqslant\epsilon_{0}|\xi|$ holds for any $1\leqslant j<k\leqslant N$. In the following Proposition 2.2 and proof of Proposition 2.3, we use $\omega_{0}\in\mathbb{S}^{d(N-1)-1}\setminus(\cup_{\leqslant j<k\leqslant N}\ker L_{jk})$, $\epsilon_{0}>0$, and $\delta>0$ chosen above.
Proposition 2.2.
Let $\phi\in\mathscr{S}(\mathbb{R}^{d(N-1)})$ be such that $\mathscr{F}_{d(N-1)}\phi\in C_{0}^{\infty}(\mathbb{R}^{d(N-1)})$ with $\operatorname{supp}\mathscr{F}_{d(N-1)}\phi\subset\{\xi\in\mathscr{C}_{\omega_{0},\delta}\bigm||\xi|\geqslant\epsilon\}$ for $0<\epsilon\leqslant\epsilon_{0}$, where $\mathscr{F}_{d(N-1)}$ denotes the Fourier transform on $L^{2}(\mathbb{R}^{d(N-1)})$. Then
$\|F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|$
$\lesssim_{\nu}t^{(d/2-\nu)(1-2\lambda)}\|\langle r_{k}-r_{j}\rangle^{\nu}\phi\|$ (2.11)
holds for any $\nu\in\mathbb{N}$, $t\geqslant T$ and $1\leqslant j<k\leqslant N$, where $\mu_{jk}=m_{j}m_{k}/(m_{j}+m_{k})$ and $\lesssim_{\nu}$ means that the constant depends on $\nu$.
Proof of Proposition 2.2.
We have
$F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi$
$=e^{-{\rm i}t^{1-2\lambda}\sum_{j=2}^{N-1}D_{j}^{2}/(2(1-2\lambda))}F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\phi,$ (2.12)
where we reconstruct the Jacobi coordinates by relabeling them so that the relative coordinate $y_{1}=r_{k}-r_{j}$ appears as the first Jacobi component for $(j,k)\neq(1,2)$. Because $|\xi_{1}|=|L_{jlk}\xi|\geqslant\epsilon_{0}|\xi|\geqslant\epsilon^{2}$ holds on $\operatorname{supp}\mathscr{F}_{d(N-1)}\phi$, it suffices to prove that
$\|F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi\|_{L^{2}(\mathbb{R}^{d})}$
$\lesssim_{\nu}t^{(d/2-\nu)(1-2\lambda)}\|\langle x_{1}\rangle^{\nu}\psi\|_{L^{2}(\mathbb{R}^{d})}$ (2.13)
for $\mathscr{F}_{d}\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $\operatorname{supp}\mathscr{F}_{d}\psi\subset\{\xi\in\mathbb{R}^{d}\bigm||\xi|\geqslant\epsilon^{2}\}$ where $\mathscr{F}_{d}$ denotes the Fourier transform on $L^{2}(\mathbb{R}^{d})$. We write
$F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi$
$=\int_{\mathbb{R}^{d}}e^{{\rm i}(x_{1}\cdot\xi-t^{1-2\lambda}|\xi|^{2}/(2(1-2\lambda)))}F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))\mathscr{F}_{d}\psi(\xi){\rm d}\xi/(2\pi)^{d/2}.$ (2.14)
When $|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda)$, we have
$|x_{1}-t^{1-2\lambda}\xi/(1-2\lambda)|\geqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda)$ (2.15)
on $\operatorname{supp}\mathscr{F}_{d}\psi$. By the relation
$e^{{\rm i}(x_{1}\cdot\xi-t^{1-2\lambda}\xi^{2}/(2(1-2\lambda)))}$
$=(x_{1}-t^{1-2\lambda}\xi/(1-2\lambda))\cdot(-{\rm i}\nabla_{\xi})e^{{\rm i}(x_{1}\cdot\xi-t^{1-2\lambda}\xi^{2}/(2(1-2\lambda)))}/|x_{1}-t^{1-2\lambda}\xi/(1-2\lambda)|^{2}$ (2.16)
and integrating by parts, we have
$|F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi(x_{1})|$
$\lesssim t^{-\nu(1-2\lambda)}F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))\|\langle x\rangle^{\nu}\psi\|_{L^{2}(\mathbb{R}_{x}^{d})}.$ (2.17)
Integrating (2.17) with respect to $x_{1}$ yields (2.13). ∎
Proposition 2.3.
The wave operators (1.20) exist.
Proof of Proposition 2.3.
We prove the existence of (2.9) for $t\rightarrow\infty$. The derivative at $t$ of $U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}$ is
$({\rm d}/{\rm d}t)U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}$
$={\rm i}\sum_{1\leqslant j<k\leqslant N}U_{\lambda}(t,T)^{*}V_{jk}(t^{\lambda}(r_{k}-r_{j}))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}.$ (2.18)
By Assumption 1.1,
$\|V_{jk}^{\rm bdd}(t^{\lambda}(r_{k}-r_{j}))F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|>\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))\|\lesssim t^{-\rho(1-\lambda)}$ (2.19)
holds for any $0<\epsilon\leqslant\epsilon_{0}$. Whereas we have
$V_{jk}^{\rm sing}(t^{\lambda}(r_{k}-r_{j}))F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|>\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))\langle D/t^{\lambda}\rangle^{-2}=0$ (2.20)
for $t\gg 1$ because $V_{jk}^{\rm sing}$ is compactly supported. By Proposition 2.2, we clearly have
$\|V_{jk}^{\rm bdd}(t^{\lambda}(r_{k}-r_{j}))F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|$
$\lesssim t^{(d/2-\nu)(1-2\lambda)}\|\langle r_{k}-r_{j}\rangle^{\nu}\phi\|$ (2.21)
because $V_{jk}^{\rm bdd}$ is bounded. Let $\chi\in C^{\infty}(\mathbb{R}^{d})$ such that
$\chi(x)=\begin{cases}\ 1&\mbox{if}\quad|x|\leqslant 2\\ \ 0&\mbox{if}\quad|x|\geqslant 3.\end{cases}$ (2.22)
Reconstructing the Jacobi coordinates for $(j,k)\not=(1,2)$, we have
$\|V_{jk}^{\rm sing}(t^{\lambda}(r_{k}-r_{j}))F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|$
$\leqslant\|V_{jk}^{\rm sing}(t^{\lambda}x_{1}/\sqrt{\mu_{jk}})\chi((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\phi\|.$ (2.23)
Noting that $\|V_{jk}^{\rm sing}(t^{\lambda}x_{1}/\sqrt{\mu_{jk}})\langle\sqrt{\mu_{jk}}D_{1}/t^{\lambda}\rangle^{-2}\|=\|V_{jk}^{\rm sing}(x_{1})\langle D_{1}\rangle^{-2}\|$ and that
${\rm i}[D_{1}^{2},\chi((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))]=((1-2\lambda)/\epsilon^{2}t^{1-2\lambda})(\nabla\chi)((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))\cdot D_{1}$
$-{\rm i}((1-2\lambda)/\epsilon^{2}t^{1-2\lambda})^{2}(\Delta\chi)((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda})),$ (2.24)
we have
$\|\langle\sqrt{\mu_{jk}}D_{1}/t^{\lambda}\rangle^{2}\chi((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\phi\|$
$\lesssim\|\chi((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\phi\|$
$+t^{-1}\|(\nabla\chi)((1-2\lambda)x_{1}/(\epsilon^{2}t^{1-2\lambda}))\cdot e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}D_{1}\phi\|+t^{-2+2\lambda}\|\phi\|.$ (2.25)
Choosing $\epsilon>0$ sufficiently small, we can use Proposition 2.2 again for the first and second terms on the right side of (2.25). By (2.19), (2.20), (2.21), and (2.25). We therefore have
$\|({\rm d}/{\rm d}t)U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|\lesssim t^{-\rho(1-\lambda)}+t^{(d/2-\nu)(1-2\lambda)}+t^{-2+2\lambda}$ (2.26)
for $t\gg 1$. We note that $\rho(1-\lambda)>1$, $2-2\lambda>1$, and $(\nu-d/2)(1-2\lambda)>1$ for $\nu\gg 1$. By the Cook-Kuroda method ([45, Theorem XI.4]), this completes the proof. ∎
3 Inverse Scattering
To prove Theorem 1.2, it is convenient to choose different Jacobi coordinates from those of (1.6). According to [16, 49, 53], we focus on the pair of particles $1$ and $2$ and prove that $V_{12,1}=V_{12,2}$ because the proofs for the other pairs for $j$ and $k$ can be demonstrated in the same analogy. Let $y_{12}=r_{2}-r_{1}$ be the relative position between $r_{1}$ and $r_{2}$, and
$y_{j}=r_{j}-(m_{1}r_{1}+m_{2}r_{2})/(m_{1}+m_{2})$ (3.1)
for $3\leqslant j\leqslant N$ be the relative position between $r_{j}$ and the center of mass for $r_{1}$ and $r_{2}$. Their conjugate momentums are given by
${\rm i}\nabla_{y_{12}}=\mu_{12}(-{\rm i}\nabla_{r_{2}}/m_{2}-(-{\rm i}\nabla_{r_{1}})/m_{1})$ (3.2)
and
${\rm i}\nabla_{y_{j}}=\mu_{j}(-{\rm i}\nabla_{r_{j}}/m_{j}-(-{\rm i}\nabla_{r_{1}}-{\rm i}\nabla_{r_{2}})/(m_{1}+m_{2}))$ (3.3)
for $3\leqslant j\leqslant N$ where
$\mu_{12}=m_{1}m_{2}/(m_{1}+m_{2})$ (3.4)
and
$\mu_{j}=m_{j}(m_{1}+m_{2})/(m_{j}+m_{1}+m_{2})$ (3.5)
for $3\leqslant j\leqslant N$ are the reduced masses respectively. Because of the relations
$\sum_{j=1}^{N}m_{j}r_{j}^{2}=Mr_{\rm cm}^{2}+\mu_{12}y_{12}^{2}+\sum_{j=3}^{N}\mu_{j}y_{j}^{2},$ (3.6)
$\sum_{j=1}^{N}\Delta_{r_{j}}/m_{j}=\Delta_{\rm cm}/M+\Delta_{y_{12}}/\mu_{12}+\sum_{j=3}^{N}\Delta_{y_{j}}/\mu_{j},$ (3.7)
we rewrite the free Hamiltonian $H_{0}(t)$ in (2.1) using $D=(D_{12},D_{3},\ldots,D_{N})$ and $x=(x_{12},x_{3}\ldots,x_{N})$ with $x_{12}=\sqrt{\mu_{12}}y_{12}$ and $x_{j}=\sqrt{\mu_{j}}y_{j}$ for $3\leqslant j\leqslant N$. The relative momentum $D_{jk}$ is defined such that $D_{1k}=D_{2k}=D_{k}-D_{12}$ for $3\leqslant k\leqslant N$ and $D_{jk}=D_{k}-D_{j}$ for $3\leqslant j<k\leqslant N$. These coordinates are different from those in [16, 49, 53] because of the scalings of $x_{12}$ and $x_{j}$ for $3\leqslant j\leqslant N$. By virtue of these scalings, we omit the effects of reduced masses in our notations. Even in these coordinates, the wave operators (1.20) exist by the same proofs in Section 2. In particular, the strong limits
$W_{\lambda}^{\pm}=\operatorname*{s-lim}_{t\rightarrow\pm\infty}U(t,0)^{*}U_{0,\lambda}(t)$ (3.8)
also exist, and we define
$S_{\lambda}(V)=(W_{\lambda}^{+})^{*}W_{\lambda}^{-}.$ (3.9)
Because $W^{\pm}=W_{\lambda}^{\pm}U_{0,\lambda}(s_{\pm})^{*}U_{0}(s_{\pm},0)$ holds for $s_{+}\geqslant T$ and $s_{-}\leqslant-T$, we have
$S(V)=U_{0}(s_{+},0)^{*}U_{0,\lambda}(s_{+})S_{\lambda}(V)U_{0,\lambda}(s_{-})^{*}U_{0}(s_{-},0).$ (3.10)
This implies that $S(V_{1})=S(V_{2})$ is equivalent to $S_{\lambda}(V_{1})=S_{\lambda}(V_{2})$.
$H_{0}(t)\equiv H_{0}=D^{2}/2+\omega^{2}x^{2}/2$ is the time-independent harmonic oscillator for $|t|<T$, and its time-evolution $e^{-{\rm i}tH_{0}}$ is governed by the Mehler formula in [12, Section 2.2] and [41, Theorem 5.29], which is
$e^{-{\rm i}tH_{0}}=\mathscr{M}_{d(N-1)}(\tan\omega t/\omega)\mathscr{D}_{d(N-1)}(\sin\omega t/\omega)\mathscr{F}_{d(N-1)}\mathscr{M}(\tan\omega t/\omega)$ (3.11)
for $\omega t\notin\pi\mathbb{Z}$, where $\mathscr{M}_{d(N-1)}$ and $\mathscr{D}_{d(N-1)}$ are multiplication and dilation given by
$\mathscr{M}_{d(N-1)}(t)\phi(x)=e^{{\rm i}x^{2}/(2t)}\phi(x),$ (3.12)
$\mathscr{D}_{d(N-1)}(t)\phi(x)=({\rm i}t)^{-d(N-1)/2}\phi(x/t).$ (3.13)
Moreover,
$e^{-{\rm i}tH_{0}}={\rm i}^{d(N-1)/2}\mathscr{M}_{d(N-1)}(-\cot\omega t/\omega)\mathscr{D}_{d(N-1)}(\cos\omega t)e^{-{\rm i}\tan\omega tD^{2}/(2\omega)}$ (3.14)
for $\omega t\notin(\pi/2)\mathbb{Z}$.
For $v\in\mathbb{R}^{d}$, the normalization is $\hat{v}=v/|v|$. Let $\hat{v},e_{3},\ldots,e_{N}\in\mathbb{R}^{d}$ be unit vectors pointing in mutually different directions. We define $v_{jk}=v_{k}-v_{j}$ for $1\leqslant j<k\leqslant N$ where $v_{1}=-v/2$, $v_{2}=v/2$, and $v_{j}=|v|^{2}e_{j}$ for $3\leqslant j\leqslant N$. Let $\Phi_{0}\in\mathscr{S}(\mathbb{R}^{d(N-1)})$ be $\Phi_{0}=\phi_{1}\otimes\phi_{2}$ such that $\mathscr{F}_{d}\phi_{1}\in C_{0}^{\infty}(\mathbb{R}^{d})$ and $\mathscr{F}_{d(N-2)}\phi_{2}\in C_{0}^{\infty}(\mathbb{R}^{d(N-2)})$ with $\|\phi_{2}\|_{L^{2}(\mathbb{R}^{d(N-2)})}=1$. In addition, we assume that $\operatorname{supp}\mathscr{F}_{d}\phi_{1}\subset\{\xi_{12}\in\mathbb{R}^{d}\bigm||\xi_{12}|<\eta_{12}\}$ for some $\eta_{12}>0$ and $\operatorname{supp}\mathscr{F}_{d(N-2)}\phi_{2}\subset\{(\xi_{3},\ldots,\xi_{N})\in\mathbb{R}^{d(N-2)}\bigm||\xi_{3}|<1,\ldots,|\xi_{N}|<1\}$. We define $\Phi_{v}=\mathscr{T}_{v}\Phi_{0}$ for
$\mathscr{T}_{v}=e^{{\rm i}v\cdot x_{12}}\prod_{j=3}^{N}e^{{\rm i}v_{j}\cdot x_{j}}.$ (3.15)
Then there exists $f_{jk}\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $\operatorname{supp}f_{jk}\subset\{\xi\in\mathbb{R}^{d}\bigm||\xi|\leqslant\eta_{jk}\}$ where $\eta_{1k}=\eta_{2k}=1+\eta_{12}$ for $3\leqslant k\leqslant N$ and $\eta_{jk}=2$ for $3\leqslant j<k\leqslant N$ such that
$\Phi_{v}=\mathscr{T}_{v}f_{jk}(D_{jk})\Phi_{0}=f_{jk}(D_{jk}-v_{jk})\Phi_{v}.$ (3.16)
To apply the Enss-Weder time-dependent method [16], the following reconstruction formula is essential. By virtue of this formula, the uniqueness of the potential functions follows from the injective property of the Radon transform.
Theorem 3.1.
Let $\Phi_{v}$ be defined as (3.16) and $\Psi_{v}$ have the same properties for $\Psi_{0}=\psi_{1}\otimes\phi_{2}$. Then
$\lim_{|v|\rightarrow\infty}|v|({\rm i}(S_{\lambda}(V)-1)\Phi_{v},\Psi_{v})=\int_{-\infty}^{\infty}(V_{12}(r_{2}-r_{1}+\hat{v}t)\phi_{1},\psi_{1})_{L^{2}(\mathbb{R}^{d})}{\rm d}t$ (3.17)
holds for $V=\sum_{1\leqslant j<k\leqslant k}V_{jk}$, which satisfies Assumption 1.1.
It seems difficult to study the time evolution of $U_{0}(t,0)$ directly for all $t\in\mathbb{R}$. Instead, we analyze $e^{-{\rm i}tH_{0}}$ for $|t|<T$ and $e^{\mp{\rm i}|t|^{1-2\lambda}D^{2}/(2(1-2\lambda))}$ for $|t|\geqslant T$ according to [26]. We define $U_{0,\lambda}(t)=e^{-{\rm i}tH_{0}}$ for $|t|<T$. In the proofs below, we can assume that
$\pi/(2\omega)\leqslant T<\pi/\omega$ (3.18)
without loss of generality, as in [26].
Lemma 3.2.
Let $\Phi_{v}$ be as in Theorem 3.1. Then
$\int_{-\infty}^{\infty}\|V_{jk}^{\rm bdd}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|{\rm d}t=O(|v_{jk}|^{-1})$ (3.19)
holds as $|v|\rightarrow\infty$ for any $1\leqslant j<k\leqslant N$.
Proof of Lemma 3.2.
We divide the integral such that
$\int_{-\infty}^{\infty}=\int_{|t|<T}+\int_{|t|\geqslant T}$ (3.20)
and first consider the integral on $|t|<T$. By (3.14) and (3.16), we have
$\|V_{jk}^{\rm bdd}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|=\|V_{jk}^{\rm bdd}(\cos\omega t(r_{k}-r_{j}))e^{-{\rm i}\tan\omega tD^{2}/(2\omega)}f_{jk}(D_{jk}-v_{jk})\Phi_{v}\|$
$=\|V_{jk}^{\rm bdd}(\cos\omega tx_{12}/\sqrt{\mu_{jk}})e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}f_{jk}(D_{12}-v_{jk})\Phi_{v}\|$
$=\|V_{jk}^{\rm bdd}((\cos\omega tx_{12}+\sin\omega tv_{jk}/\omega)/\sqrt{\mu_{jk}})e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}f_{jk}(D_{12})e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|,$ (3.21)
where we have reconstructed the Jacobi coordinates from $y_{12}=r_{k}-r_{j}$ for $(j,k)\not=(1,2)$ and used the relation
$e^{-{\rm i}v_{jk}\cdot x_{12}}e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}e^{{\rm i}v_{jk}\cdot x_{12}}=e^{-{\rm i}\tan\omega t|v_{jk}|^{2}/(2\omega)}e^{-{\rm i}\tan\omega tD_{12}\cdot v_{jk}/\omega}e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}.$ (3.22)
From [26, (39), (40), and (43) in the proof of Lemma 2.3], we have
$\int_{|t|<T}\|V_{jk}^{\rm bdd}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|{\rm d}t=O(|v_{jk}|^{-1}).$ (3.23)
We next consider the integral on $|t|\geqslant T$. In particular, we consider $t\geqslant T$. By (2.2) and (3.16), we have
$\|V_{jk}^{\rm bdd}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|=\|V_{jk}^{\rm bdd}(t^{\lambda}(r_{k}-r_{j}))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}f_{jk}(D_{jk}-v_{jk})\Phi_{v}\|$
$=\|V_{jk}^{\rm bdd}(t^{\lambda}x_{12}/\sqrt{\mu_{jk}})e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}f_{jk}(D_{12}-v_{jk})\Phi_{v}\|$
$=\|V_{jk}^{\rm bdd}((t^{\lambda}x_{12}+t^{1-2\lambda}v_{jk}/(1-2\lambda))/\sqrt{\mu_{jk}})e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}f_{jk}(D_{12})e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|,$ (3.24)
where we have reconstructed the Jacobi coordinates from $y_{12}=r_{k}-r_{j}$ for $(j,k)\not=(1,2)$ and used the relation
$e^{-{\rm i}v_{jk}\cdot x_{12}}e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}e^{{\rm i}v_{jk}\cdot x_{12}}$
$=e^{-{\rm i}t^{1-2\lambda}|v_{jk}|^{2}/(2(1-2\lambda))}e^{-{\rm i}t^{1-2\lambda}D_{12}\cdot v_{jk}/(1-2\lambda)}e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}.$ (3.25)
By [26, (47) and (51) in the proof of Lemma 2.3], we have
$\int_{|t|\geqslant T}\|V_{jk}^{\rm bdd}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|{\rm d}t=O(|v_{jk}|^{-1}).$ (3.26)
(3.23) and (3.26) complete the proof. ∎
Lemma 3.3.
Let $\Phi_{v}$ be as in Theorem 3.1. Then
$\int_{-\infty}^{\infty}\|V_{jk}^{\rm sing}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|{\rm d}t=O(|v_{jk}|^{-1})$ (3.27)
holds as $|v|\rightarrow\infty$ for any $1\leqslant j<k\leqslant N$.
Proof of Lemma 3.3.
We divide the integral such that
$\int_{-\infty}^{\infty}=\int_{|t|\leqslant\pi/(4\omega)}+\int_{\pi/(4\omega)<|t|<T}+\int_{|t|\geqslant T}$ (3.28)
and first consider the integral on $|t|\leqslant\pi/(4\omega)$. As we similarly had for (3.21), we have
$\|V_{jk}^{\rm sing}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|=\|V_{jk}^{\rm sing}((\cos\omega tx_{12}+\sin\omega tv_{jk}/\omega)/\sqrt{\mu_{jk}})\langle D_{12}/\cos\omega t\rangle^{-2}$
$\times e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}f_{jk}(D_{12})\langle D_{12}/\cos\omega t\rangle^{2}e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|.$ (3.29)
We take $g_{jk}\in C_{0}^{\infty}(\mathbb{R}^{d})$ such that $f_{jk}=f_{jk}g_{fk}$. Noting that
$\|\langle x_{12}\rangle^{2}\langle D_{12}/\cos\omega t\rangle^{2}g_{jk}(D_{12})e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|$
$\lesssim\|\langle D_{12}\rangle^{2}g_{jk}(D_{12})\|\|\langle x_{12}\rangle^{2}\Phi_{0}\|+\|[x_{12}^{2},\langle D_{12}\rangle^{2}g_{jk}(D_{12})]e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|\lesssim 1,$ (3.30)
we have
$\int_{|t|\leqslant\pi/(4\omega)}\|V_{jk}^{\rm sing}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|=O(|v_{jk}|^{-1})$ (3.31)
by [26, (74) and (79) in the proof of Lemma 3.2]. We next consider the integral on $\pi/(4\omega)<|t|<T$. We define the $d$-dimensional harmonic oscillator
$H_{0,12}=D_{12}^{2}/2+w^{2}x_{12}^{2}/2$ (3.32)
and note that the Mehler formula
$e^{-{\rm i}H_{0,12}}={\rm i}^{d/2}\mathscr{M}_{d}(-\cot\omega t/\omega)\mathscr{D}_{d}(\cos\omega t)e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}$ (3.33)
also holds. As we similarly had for (3.21), we have
$\|V_{jk}^{\rm sing}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|$
$=\|V_{jk}^{\rm sing}((\cos\omega tx_{12}+\sin\omega tv_{jk}/\omega)/\sqrt{\mu_{jk}})e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|$
$=\|V^{\rm sing}((x_{12}+\sin\omega tv_{jk}/\omega)/\sqrt{\mu_{jk}})e^{-{\rm i}H_{0,12}}e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|$ (3.34)
using (3.33). We therefore have
$\int_{\pi/(4\omega)<|t|<T}\|V_{jk}^{\rm sing}(r_{k}-r_{j})e^{-{\rm i}tH_{0}}\Phi_{v}\|=O(|v_{jk}|^{-1})$ (3.35)
by [26, (88) and (93) in the proof of Lemma 3.2]. Finally, we consider the integral on $|t|\geqslant T$, in particular $t\geqslant T$. As we similarly had for (3.24), we have
$\|V_{jk}^{\rm sing}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|=\|V_{jk}^{\rm sing}((t^{\lambda}x_{12}+t^{1-2\lambda}v_{jk}/(1-2\lambda))/\sqrt{\mu_{jk}})\langle D_{12}/t^{\lambda}\rangle^{-2}$
$\times e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}f_{jk}(D_{12})\langle D_{12}/t^{\lambda}\rangle^{2}e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|.$ (3.36)
As in (3.30), we have
$\|\langle x_{12}\rangle^{2}\langle D_{12}/t^{\lambda}\rangle^{2}g_{jk}(D_{12})e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|\lesssim 1$ (3.37)
for $g_{jk}\in C_{0}^{\infty}(\mathbb{R}^{d})$ such that $f_{jk}=f_{jk}g_{jk}$. We inductively have
$\|\langle x_{12}\rangle^{2}\langle D_{12}/t^{\lambda}\rangle^{N}g_{jk}(D_{12})e^{-{\rm i}v_{jk}\cdot x_{12}}\Phi_{v}\|\lesssim 1$ (3.38)
for any $N\in\mathbb{N}$. Therefore, by [26, (97) and (99) in the proof of Lemma 3.2], we have
$\int_{|t|\geqslant T}\|V_{jk}^{\rm sing}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|=O(|v_{jk}|^{-1}).$ (3.39)
(3.31), (3.35), and (3.39) imply that Lemma 3.3 holds. ∎
By virtue of Lemmas 3.2 and 3.3, we immediately have the following Lemma as in [16, Corollary 2.3] (see also [4, 6, 7, 10, 11, 21, 22, 23, 26, 27, 42, 43, 44, 49, 53]). We therefore omit its proof.
Lemma 3.4.
Let $\Phi_{v}$ be as in Theorem 3.1. Then
$\sup_{t\in\mathbb{R}}\|(U(t,0)W_{\lambda}^{-}-U_{0,\lambda}(t))\Phi_{v}\|=O(|v|^{-1})$ (3.40)
holds as $|v|\rightarrow\infty$.
Proof of Theorem 3.1.
We have
${\rm i}(S_{\lambda}(V)-1)={\rm i}(W_{\lambda}^{+}-W_{\lambda}^{-})^{*}W_{\lambda}^{-}={\rm i}\int_{-\infty}^{\infty}(({\rm d}/{\rm d}t)U(t,0)^{*}U_{0,\lambda}(t))^{*}W_{\lambda}^{-}{\rm d}t$
$=\sum_{1\leqslant j<k\leqslant N}\int_{-\infty}^{\infty}U_{0,\lambda}(t)^{*}V_{jk}(r_{k}-r_{j})U(t,0)W_{\lambda}^{-}{\rm d}t$ (3.41)
and
$|v|(({\rm i}S(V)-1)\Phi_{v},\Psi_{v})=|v|\int_{-\infty}^{\infty}(V_{12}(r_{2}-r_{1})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v}){\rm d}t+R(v),$ (3.42)
where
$R(v)=|v|\sum_{(j,k)\not=(1,2)}\int_{-\infty}^{\infty}(V_{jk}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v}){\rm d}t$
$+|v|\sum_{1\leqslant j<k\leqslant N}\int_{-\infty}^{\infty}((U(t,0)W_{\lambda}^{-}-U_{0,\lambda}(t))\Phi_{v},V_{jk}(r_{k}-r_{j})U_{0,\lambda}(t)){\rm d}t.$ (3.43)
Noting that
$\int_{-\infty}^{\infty}\|V_{jk}(r_{k}-r_{j})U_{0,\lambda}(t)\Phi_{v}\|{\rm d}t=O(|v|^{-2})$ (3.44)
holds as $|v|\rightarrow\infty$ for $(j,k)\not=(1,2)$, we have $R(v)=O(|v|^{-1})$ by Lemmas 3.2, 3.3, and 3.4. We now prove
$|v|\int_{-\infty}^{\infty}(V_{12}(r_{2}-r_{1})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v}){\rm d}t\rightarrow\int_{-\infty}^{\infty}(V_{12}(r_{2}-r_{1}+\hat{v}t)\phi_{1},\psi_{1})_{L^{2}(\mathbb{R}^{d})}{\rm d}t$ (3.45)
as $|v|\rightarrow\infty$. We first focus on the term $V_{12}^{\rm bdd}$. We divide the integral such that
$\int_{-\infty}^{\infty}=\int_{|t|<\pi/(2\omega)}+\int_{\pi/(2\omega)\leqslant|t|<T}+\int_{|t|\geqslant T}$ (3.46)
and consider the integrals on $|t|<\pi/(2\omega)$ and $\pi/(2\omega)\leqslant|t|<T$. Recalling that $\phi_{2}$ is normalized, we have
$(V_{12}^{\rm bdd}(r_{2}-r_{1})e^{-{\rm i}tH_{0}}\Phi_{v},e^{-{\rm i}tH_{0}}\Psi_{v})$
$=(V_{12}^{\rm bdd}(\cos\omega tx_{12}/\sqrt{\mu_{12}})e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}\Phi_{v},e^{-{\rm i}\tan\omega tD_{12}^{2}/(2\omega)}\Psi_{v})$
$=(V_{12}^{\rm bdd}(x_{12}/\sqrt{\mu_{12}})e^{-{\rm i}H_{0,12}}e^{{\rm i}v\cdot x_{12}}\phi_{1},e^{-{\rm i}H_{0,12}}e^{{\rm i}v\cdot x_{12}}\psi_{1})_{L^{2}(\mathbb{R}^{d})}$ (3.47)
for $|t|<T$ by (3.14) and (3.33). Therefore
$|v|\int_{|t|<\pi/(2\omega)}(V_{12}^{\rm bdd}(r_{2}-r_{1})e^{-{\rm i}tH_{0}}\Phi_{v},e^{-{\rm i}tH_{0}}\Psi_{v}){\rm d}t$
$\rightarrow\int_{-\infty}^{\infty}(V_{12}^{\rm bdd}(x_{12}/\sqrt{\mu_{12}}+\hat{v}t)\phi_{1},\psi_{1})_{L^{2}(\mathbb{R}^{d})}{\rm d}t$ (3.48)
and
$|v|\int_{\pi/(2\omega)\leqslant|t|<T}(V_{12}^{\rm bdd}(r_{2}-r_{1})e^{-{\rm i}tH_{0}}\Phi_{v},e^{-{\rm i}tH_{0}}\Psi_{v}){\rm d}t\rightarrow 0$ (3.49)
hold as $|v|\rightarrow\infty$ by [26, (64) and (65) in the proof of Theorem 2.1]. Let $U_{0,12}(t,s)$ be the propagator for the $d$-dimensional time-dependent harmonic oscillator
$H_{0,12}(t)=D_{12}^{2}/2+k(t)x_{12}^{2}/2$ (3.50)
and define
$U_{0,12,\lambda}(t)=e^{{\rm i}\lambda x_{12}^{2}/(2t)}e^{-{\rm i}\lambda\log tA_{12}}e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}$ (3.51)
if $t\geqslant T$ and
$U_{0,12,\lambda}(t)=e^{{\rm i}\lambda x_{12}^{2}/(2t)}e^{-{\rm i}\lambda\log(-t)A_{12}}e^{{\rm i}(-t)^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}$ (3.52)
if $t\leqslant-T$, where $A_{12}=(D_{12}\cdot x_{12}+x_{12}\cdot D_{12})/2$. By (2.2), (2.3), (3.51), and (3.52), we have
$(V_{12}^{\rm bdd}(r_{2}-r_{1})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v})$
$=(V_{12}^{\rm bdd}(t^{\lambda}x_{12}/\sqrt{\mu_{12}})e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}\Phi_{v},e^{-{\rm i}t^{1-2\lambda}D_{12}^{2}/(2(1-2\lambda))}\Psi_{v})$
$=(V_{12}^{\rm bdd}(x_{12}/\sqrt{\mu_{12}})U_{0,12,\lambda}(t)e^{{\rm i}v\cdot x_{12}}\phi_{1},U_{0,12,\lambda}(t)e^{{\rm i}v\cdot x_{12}}\psi_{1})_{L^{2}(\mathbb{R}^{d})}$ (3.53)
for $|t|\geqslant T$. By [26, (66) in the proof of Theorem 2.1],
$|v|\int_{|t|\geqslant T}(V_{12}^{\rm bdd}(r_{2}-r_{1})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v}){\rm d}t\rightarrow 0$ (3.54)
holds as $|v|\rightarrow\infty$. We next focus on the term $V_{12}^{\rm sing}$. We divide the integral such that
$\int_{-\infty}^{\infty}=\int_{|t|<\pi/(4\omega)}+\int_{\pi/(4\omega)\leqslant|t|<T}+\int_{|t|\geqslant T}$ (3.55)
and consider the integrals on $|t|<\pi/(4\omega)$ and $\pi/(4\omega)\leqslant|t|<T$. Because (3.47) also holds with $V_{12}^{\rm bdd}$ replaced by $V_{12}^{\rm sing}$, we have
$|v|\int_{|t|<\pi/(4\omega)}(V_{12}^{\rm sing}(r_{2}-r_{1})e^{-{\rm i}tH_{0}}\Phi_{v},e^{-{\rm i}tH_{0}}\Psi_{v}){\rm d}t$
$\rightarrow\int_{-\infty}^{\infty}(V_{12}^{\rm sing}(x_{12}/\sqrt{\mu_{12}}+\hat{v}t)\phi_{1},\psi_{1})_{L^{2}(\mathbb{R}^{d})}{\rm d}t$ (3.56)
and
$|v|\int_{\pi/(4\omega)\leqslant|t|<T}(V_{12}^{\rm sing}(r_{2}-r_{1})e^{-{\rm i}tH_{0}}\Phi_{v},e^{-{\rm i}tH_{0}}\Psi_{v}){\rm d}t\rightarrow 0$ (3.57)
as $|v|\rightarrow\infty$ by [26, (110) and (111) in the proof of Theorem 3.1]. Finally, we consider the integral on $|t|\geqslant T$. Because (3.53) also holds with $V_{12}^{\rm bdd}$ replaced by $V_{12}^{\rm sing}$, we have
$|v|\int_{|t|\geqslant T}(V_{12}^{\rm bdd}(r_{2}-r_{1})U_{0,\lambda}(t)\Phi_{v},U_{0,\lambda}(t)\Psi_{v}){\rm d}t\rightarrow 0$ (3.58)
as $|v|\rightarrow\infty$ by [26, (112) in the proof of Theorem 3.1]. Combining (3.48), (3.49), (3.54), (3.56), (3.57), and (3.58), we have (3.45). ∎
Proof of Theorem 1.2.
We assume that $S_{\lambda}(V_{1})=S_{\lambda}(V_{2})$ because (3.10) holds. By the same computation as that of [26, (77) in the proof of Lemma 3.2], we have the condition commonly referred to as the Enss condition
$\int_{0}^{\infty}\|V_{12}(r_{2}-r_{1})\langle D_{12}\rangle^{-2}F(|r_{2}-r_{1}|\geqslant R)\|_{\mathscr{B}(L^{2}(\mathbb{R}^{d}))}{\rm d}R<\infty.$ (3.59)
By virtue of Theorem 3.1 and the Plancherel formula for the Radon transform ([19, Theorem 2.17 in Chapter 1]), we have $V_{12,1}=V_{12,2}$ in the same way as for the proof of [16, Theorem 1.1]. ∎
Acknowledgments. This work was supported by JSPS KAKENHI Grant Number JP21K03279.